solve -m squared= -81
(i couldnt figure out how to make the square symbol)

Answer by stanbon(75887)

solve -m squared= -81
--------------
m^2 - 81 = 0
(m-9)(m+9) = 0
m=9 or m=-9
================
Cheers,
Stan H.